OpenStudy (anonymous):

how do I get 3sqrt(8)+3sqrt(8)=32^(5/2) ?

OpenStudy (anonymous):

\[3\sqrt{8}+3\sqrt{8}=32^{\frac{ 5 }{ 2 }}\]

terenzreignz (terenzreignz):

At face-value, this doesn't seem to be true... \[\large 2 < \sqrt{8} < 3\]\[\large 6 < 3\sqrt{8} < 9\]

OpenStudy (anonymous):

it comes from \[3\sqrt{ 9t2−7t+8 }+3\sqrt{8t^2+6t+8 }\] when t=0

terenzreignz (terenzreignz):

But that above is not an equation. Do you mean that expression above is always equal to \[\huge 32^{\frac52}\]?

OpenStudy (anonymous):

sorry I mean \[3\sqrt{9t2−7t+8}+3\sqrt{8t2+6t+8}=32 ^{\frac{ 5 }{ 2 }}\] when t=0

terenzreignz (terenzreignz):

Well, then, this does indeed, lead to your original equation, but I still don't see any truth in it :D If this equation is to be true, it must be for some other t...

OpenStudy (anonymous):

ok so it must be a mistake in the solutions. I think it should say 3*2^(5/2). so.. sqrt{8}=2*sqrt{2}=2^(3/2) and 2*2*sqrt{2}=2^(5/2) great

terenzreignz (terenzreignz):

Oh, so that was it...